A monochromatic beam of wavelength 6000A and intensity 10mwcm-2 fall on surface the no of photons striking the surface per second per cm2 are

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
I = (n/tA )*(hc/Lambda)........1
no of electrons emitted = 1/100 * (nt/A)
Intensity = 10
Put all values in equaton 1 
(n/tA) = 3.01 * 10^19 
So, no of photons incident per unit time per unit area 
= 0.01 8 (n/tA) 
= 3 * 10^17 
Hope this helps